# Modal Verification Patterns for Systems

### v1

@@ 3.2 Resource
When considering weak memory models, we also want points-to information (address-value mappings).
@@ 3.3 Nominalization
This choice trades off complexity against flexibility and scoping constraints. Strongly nominalized elements track additional specifier/prover-visible book-keeping data. But in exchange for carrying those identifiers of their context with them, strongly nominalized elements can be used together under any modality. For example, one can use the StackRegion modality to talk about two different stack frames simultaneously for code which accesses multiple stack frames: $\framebox{n}\;\ell\mapsto v\ast\framebox{n+1}\;\ell^{\prime}\mapsto v^{\prime}$. Using a given strongly nominalized assertion element under different modalities for different frames does not change its meaning.
@@ 4 Conclusion
We have argued the essential patterns that help to choose the modalities when working out how to specify different kinds of systems code based on recent successes. It also captures (a slightly more organized version of) our own thinking in coming up with designs for specifying distributed systems [12], virtual memory managers [19], and in ongoing work using different modalities for different parts of a copy-on-write filesystem (e.g., for assertions true in different snapshots). So not only are the modalities useful for conducting proofs, the pieces we have identified seem to be an effective way of working out a specific modality for a specific use-case.
As our main conclusion and future work, we think that these essential patterns in the designs of modal abstractions discussed constitute the fundamentals of verification patterns when working out how to specify different kinds of systems code.

### v2

@@ 3.2 Resource
When considering weak memory models, we consider address-value mappings of thread-local views in C11 memory model [34]). Modalities $\Delta_{\pi}(P)$ and $\nabla_{\pi}(P)$ with $P$ represent the memory address-value mappings that held before or will hold after certain memory fence operations by thread $\pi$.
@@ 3.3 Nominals for Systems Resource Contexts
This choice trades off complexity against flexibility and scoping constraints. Strongly nominalized elements track additional specifier/prover-visible book-keeping data. But in exchange for carrying those identifiers of their context with them, strongly nominalized elements can be used together under any modality. For example, one can use the StackRegion modality to talk about two different stack frames simultaneously for code which accesses multiple stack frames: $\framebox{n}\;\ell\mapsto v\ast\framebox{n+1}\;\ell^{\prime}\mapsto v^{\prime}$. Using a given strongly nominalized assertion element under different modalities for different frames does not change its meaning. Likewise, weak-memory specific modalities existentially quantify over other views, related to the “current” view (the one where the current thread’s assertions are evaluated), and evaluate a resource (a thread-local view of an address-value mapping) with respect to those other views. Given that views are identified by nominals, referring to all the other views makes weak-memory modalities strongly nominalized.
@@ 4 Conclusion
We have argued that there is a systematic pattern emerging in the use of modalities in the verification of low-level systems software, and explained how a range of modern work fits into this pattern. While we have focused our attention on public verification work carried out in Iris, it also captures (a slightly more organized version of) our own thinking in coming up with designs for specifying distributed systems [19], and in ongoing work using different modalities for different parts of a copy-on-write filesystem (e.g., for assertions true in different snapshots). So not only are the modalities observably present, but they are demonstrably useful for conducting proofs of systems software. We believe that these essential patterns in the designs of modal abstractions for systems verification constitute a fundamental concept when working out how to specify different kinds of systems code.